b) Test the series for convergence or divergence. 00 (-3)*+ Σ k = 1 42%
Answer #1

solution! - ktl -3) 216 z K=1 2 (-3) & (-3)* z k= 4 (42) k (-3) (-3) k (160k 2 k=1 k -3 16 16 - z -3. K=1 So, this is geomeli